A medical researcher wants to compare the pulse rates of smokers and non-smokers. He believes that the pulse rate for smokers and non-smokers is different and wants to test this claim at the 0.050.05 level of significance. A sample of 3333 smokers has a mean pulse rate of 9090, and a sample of 5050 non-smokers has a mean pulse rate of 8686. The population standard deviation of the pulse rates is known to be 55 for smokers and 66 for non-smokers. Let μ1μ1 be the true mean pulse rate for smokers and μ2μ2 be the true mean pulse rate for non-smokers.

Step 2 of 5 :  

Compute the value of the test statistic. Round your answer to two decimal places.
Answer #1

Given that,

For Smokers :

For Non-smokers :

The null and alternative hypotheses are,

H0 : μ1 = μ2

Ha : μ1 ≠ μ2

Test statistic is,

Therefore, the value of the test statistic is, Z = 3.29

p-value = 2 * P(Z > 3.21) = 2 * 0.0005 = 0.0010

=> p-value = 0.0010